OpenStudy (anonymous):

3/2(x+8) = 6x Simplify. Show your work

OpenStudy (anonymous):

1,5(x+8) = 6x 1,5x + 12 = 6x -4,5x = -12 x = -12/-4,5 = 8/3

OpenStudy (anonymous):